OpenStudy (anonymous):

Given the two functions: f(x)=3-2x and g(x)=(x-2)^2 answer the following: f(5)

OpenStudy (anonymous):

-7

OpenStudy (saifoo.khan):

f(x) = 3-10 f(x) = -7

OpenStudy (saifoo.khan):

g(x) = 3^2 = 9

OpenStudy (anonymous):

So would it be 9 or -7?

OpenStudy (saifoo.khan):

u askd about two different funtions so there will b 2 ans

OpenStudy (anonymous):

I'm doing multiple choice... both are available but can only choose one. Wierd. Thanks though!